# Function-based Labels for Complementary Recommendation: Definition, Annotation, and LLM-as-a-Judge

### v2

@@ 3.1 Definition
These relationships were intended for use in the annotation process, as described later. Labels with clearer relationship definitions were placed earlier in the sequence (A)–(E) to ensure the efficiency and reliability of annotation. Annotators use a flowchart to determine whether an item pair fits each definition in order. Based on insights from the pre-tests, we subdivided complementary relationships into finer categories (B-*, C-*) to better reflect functional distinctions and to facilitate annotation efficiency by reducing annotator burden.
@@ 3.2.2 Annotation Results
Human annotation required on average 16 JPY and 50 seconds per item pair. Of the 2,800 pairs collected through annotation, 41 that did not fit the FBLs definitions were excluded as invalid responses. This resulted in a human-annotated FBL dataset, $\mathfrak{F}_{\text{Human}}$, comprising 2,759 pairs. The label distribution in $\mathfrak{F}_{\text{Human}}$ is shown in Figure 1. Label A, which corresponds to substitute relationships, accounts for 14.9%. Labels B-* and C-*, which correspond to complementary relationships, comprised 21.4%. Label D, which represents unrelated relationships, was the most common at 48.0%. Label E, which indicates relationships that could not be defined, was relatively low at 10.8%, suggesting that the proposed function-based relationship definitions covered most relationships.
Regarding annotation reliability, the proportion of parity cases was low at 4.8%. Fleiss’ Kappa [20] further yielded an overall value of 0.63 (moderate agreement), and Figure 2 shows the Kappa values with 95% confidence intervals (95% CI) for each label. Agreement was relatively high for Labels A and B-*, whereas lower values were observed for some fine-grained complementary categories (C-3, C-4) and for Label E. These results suggest that the proposed FBLs provided sufficiently clear and consistent definitions for human annotation, while certain categories still posed greater challenges. For subsequent evaluation experiments, we used $\mathfrak{F}^{\prime}_{\text{Human}}$, which comprised 2,625 pairs, excluding the contested pairs from $\mathfrak{F}_{\text{Human}}$. The FBLs dataset developed in this study is publicly available on GitHub [^1] https://github.com/okamoto-lab/fbl_dataset. The dataset contains 2,800 item pairs with their corresponding relationship labels, including web links to each item in the pairs, to support further research on complementary recommender systems.
@@ 4.1 Settings for Machine Learning Models
As input to the traditional traditional models, we construct a 424-dimensional feature vector $\boldsymbol{x_{i}=[x^{sim}_{i},x^{mch}_{i},x^{cate}_{i}]}$, combining three types of features: (i) Similarity features: cosine similarities between multilingual-e5-large [^3] https://huggingface.co/intfloat/multilingual-e5-large embeddings of item information (title, description, specification, category), (ii) Match flag features: binary indicators of exact matches for major category, subcategory, manufacturer, and brand, and (iii) Category match features: binary indicators showing whether item pairs match in each element of the major and subcategory classifications. For the neural models, the input was constructed as a single text sequence by concatenating the attributes of each item (title, description, category, maker, and brand) in this order, separated by commas. A delimiter is then placed between the sequences of item $x$ and item $y$ to distinguish the two items.
@@ 5.3 Classification Accuracy of LLMs
Similarly, in Macro-F1, the 9-cls setting achieved higher scores than the 3-cls setting did. GPT w/ 9-cls achieved the best result with 0.849, followed by Llama w/ 9-cls at 0.822, achieving a high classification accuracy comparable to or exceeding LR (0.823) without requiring labeled data. Within the 9-cls setting, improvements in scores for complementary and substitute relationships were notable, with GPT achieving 0.797 for complementary and 0.816 for substitute relationships, exhibiting significant improvements over the 3-cls setting, with scores of 0.691 and 0.581, respectively. These results indicate that a detailed label system based on item functionality enables LLMs to grasp the relationships between items more accurately, suggesting that LLMs can function effectively as annotators mimicking human perception. Moreover, GPT-4o-mini required only 0.019 JPY and 0.67 seconds per pair, corresponding to 1/842 the cost and 1/75 the time of human annotation. It maintained Macro-F1 of 0.849, demonstrating that such accuracy can be achieved at dramatically lower cost and latency.
@@ 5.4 Comparison between Human and LLMs FBLs
Figures 4 and 5 show the confusion matrices comparing human-annotated FBLs with those generated by Llama and GPT in the 9-cls setting. In Llama, outputs for complementary relationships are concentrated on B-1 (“Item $x$ can be replenished with item yy”) indicating a relatively good classification of complementary relationships, but it struggles with distinguishing more complex relationships and their directions, such as replenishment and combination. In contrast, GPT shows a reduced tendency for localization in complementary relationships, suggesting that it can capture more complex item relationships than Llama. Both LLMs also tend to classify item pairs that humans labeled as “A. Items $x$ and $y$ have the same function and usage.” and “C-4. Combining $x$ and $y$ makes them more useful.” as “D. Items x and y have no relationship.” Examples of the former include pairs, such as envelopes of different sizes or pens with different ink properties, including mechanical pencils and oil-based pens, which are difficult for humans to judge as substitutes. Examples of the latter include pairs such as sponges and bleach or disposable chopsticks and soup cups, which can be used together in the same scene (kitchen or dining), but do not necessarily have to be used, indicating a weak relationship. Fleiss’ Kappa in human annotator agreement was also low (approximately 0.4) for C-3 and C-4, suggesting that the relationship definitions for these labels were less robust and likely contributed to the difficulty of correct labeling by LLMs. The analysis of these item pair trends suggests that LLMs tend to diverge from human judgment for item pairs with ambiguous relationships or differences, functioning as stricter annotators than humans by classifying such pairs as unrelated.
Figure 4: Correspondence between Llama-generated and human-annotated labels (color bar maximum set to 100 for visualization)
Figure 5: Correspondence between GPT-generated and human-annotated labels (color bar maximum set to 100 for visualization)

### v3

@@ 3.1 Definition
These relationships were intended for use in the annotation process, as described later. Labels with clearer relationship definitions were placed earlier in the sequence from Label A to Label E to ensure the efficiency and reliability of annotation. Annotators use a flowchart to determine whether an item pair fits each definition in order. Based on insights from the pre-tests, we subdivided complementary relationships into finer categories (Labels B-* and C-*) to better reflect functional distinctions and to facilitate annotation efficiency by reducing annotator burden.
@@ 3.2.2 Annotation Results
Human annotation required on average 16 JPY and 50 seconds per item pair. Of the 2,800 pairs collected through annotation, 41 that did not fit the FBLs definitions were excluded as invalid responses. This resulted in a human-annotated FBL dataset, $\mathfrak{F}_{\text{Human}}$, comprising 2,759 pairs. The label distribution in $\mathfrak{F}_{\text{Human}}$ is shown in Figure 1. Label A, which corresponds to substitute relationships, accounts for 14.9%. Labels B-* and C-*, which correspond to complementary relationships, comprised 21.4%. Label D, which represents unrelated relationships, was the most common at 48.0%. Label E, which indicates relationships that could not be defined, was relatively low at 10.8%, suggesting that the proposed function-based relationship definitions covered most relationships. Notably, most pairs labeled as Label E were qualitatively observed to exhibit indirect complementary through a third item (e.g., an envelope and a pen refill through a pen). Such relationships cannot be adequately represented by binary relations alone and require consideration of ternary or higher-order relations. From the viewpoint of binary relation modeling, assigning these cases to Label E can therefore be regarded as an appropriate classification.
Regarding annotation reliability, the proportion of parity cases was low at 4.8%. Fleiss’ Kappa [20] further yielded an overall value of 0.63 (moderate agreement), and Figure 2 shows the Kappa values with 95% confidence intervals (95% CI) for each label. Agreement was relatively high for Labels A and B-*, whereas lower values were observed for some fine-grained complementary categories (Labels C-3 and C-4) and for Label E. These results suggest that the proposed FBLs provided sufficiently clear and consistent definitions for human annotation, while certain categories still posed greater challenges. For subsequent evaluation experiments, we used $\mathfrak{F}^{\prime}_{\text{Human}}$, which comprised 2,625 pairs, excluding the contested pairs from $\mathfrak{F}_{\text{Human}}$. The FBLs dataset developed in this study is publicly available on GitHub [^1] https://github.com/okamoto-lab/fbl_dataset. The dataset contains 2,800 item pairs with their corresponding relationship labels, including web links to each item in the pairs, to support further research on complementary recommender systems.
@@ 4.1 Settings for Machine Learning Models
As input to the traditional models, we construct a 424-dimensional feature vector $\boldsymbol{x_{i}=[x^{sim}_{i},x^{mch}_{i},x^{cate}_{i}]}$, combining three types of features: (i) Similarity features: cosine similarities between multilingual-e5-large [^3] https://huggingface.co/intfloat/multilingual-e5-large embeddings of item information (title, description, specification, category), (ii) Match flag features: binary indicators of exact matches for major category, subcategory, manufacturer, and brand, and (iii) Category match features: binary indicators showing whether item pairs match in each element of the major and subcategory classifications. For the neural models, the input was constructed as a single text sequence by concatenating the attributes of each item (title, description, category, maker, and brand) in this order, separated by commas. A delimiter is then placed between the sequences of item $x$ and item $y$ to distinguish the two items.
@@ 5.3 Classification Accuracy of LLMs
Similarly, in Macro-F1, the 9-cls setting achieved higher scores than the 3-cls setting did. GPT w/ 9-cls achieved the best result with 0.849, followed by Llama w/ 9-cls at 0.822, achieving classification accuracy exceeding LR (0.823) without requiring labeled data. Within the 9-cls setting, improvements in scores for complementary and substitute relationships were notable, with GPT achieving 0.797 for complementary and 0.816 for substitute relationships, exhibiting significant improvements over the 3-cls setting, with scores of 0.691 and 0.581, respectively. These results indicate that a detailed label system based on item functionality enables LLMs to grasp the relationships between items more accurately, suggesting that LLMs can function effectively as annotators mimicking human perception. Moreover, GPT-4o-mini required only 0.019 JPY and 0.67 seconds per pair, corresponding to 1/842 the cost and 1/75 the time of human annotation. It maintained Macro-F1 of 0.849, demonstrating that such accuracy can be achieved at dramatically lower cost and latency.
@@ 5.4 Comparison between Human and LLMs FBLs
Figures 4 and 5 show the confusion matrices comparing human-annotated FBLs with those generated by Llama and GPT in the 9-cls setting. In Llama, outputs for complementary relationships are concentrated on Label B-1 (“Item $x$ can be replenished with item yy”) indicating a relatively good classification of complementary relationships, but it struggles with distinguishing more complex relationships and their directions, such as replenishment and combination. In contrast, GPT shows a reduced tendency for localization in complementary relationships, suggesting that it can capture more complex item relationships than Llama. Both LLMs also tend to classify item pairs that humans labeled as “A. Items $x$ and $y$ have the same function and usage.” and “C-4. Combining $x$ and $y$ makes them more useful.” as “D. Items x and y have no relationship.” Examples of the former include pairs, such as envelopes of different sizes or pens with different ink properties, including mechanical pencils and oil-based pens, which are difficult for humans to judge as substitutes. Examples of the latter include pairs such as sponges and bleach or disposable chopsticks and soup cups, which can be used together in the same scene (kitchen or dining), but do not necessarily have to be used, indicating a weak relationship. Fleiss’ Kappa in human annotator agreement was also low (approximately 0.4) for Labels C-3 and C-4, suggesting that the relationship definitions for these labels were less robust and likely contributed to the difficulty of correct labeling by LLMs. The analysis of these item pair trends suggests that LLMs tend to diverge from human judgment for item pairs with ambiguous relationships or differences, functioning as stricter annotators than humans by classifying such pairs as unrelated.
Figure 4: Correspondence between Llama-generated and human-annotated labels (counts, not percentages; color scale capped at 100 for visualization)
Figure 5: Correspondence between GPT-generated and human-annotated labels (counts, not percentages; color scale capped at 100 for visualization)